OpenStudy (anonymous):

Determine whether ΔABC ~ ΔFED. If they are similar, what postulate or theorem proves their similarity? Triangles ABC and DEF. AB measures 27. AC measures 36. BC measures 45. DE measures 20. DF measures 16. EF measures 12. Answer Angle-Angle Similarity Postulate Side-Angle-Side Similarity Postulate Side-Side-Side Similarity Postulate Similarity cannot be determined.

OpenStudy (amistre64):

if the ratio of sides is the same; id say they are similar

OpenStudy (anonymous):

SOooooo Angle-Angle Similarity Postulate ??

OpenStudy (amistre64):

abc def ab = 27 , /4 = not de = 20 , /4 = 5 bc = 45 /4 = not ef = 12 /4 = 3 ac = 36 /4 = 9 df = 16 /4 = 4 i dont think they are similar, but its been awhile since I tried this level of math

OpenStudy (amistre64):

since no angles are present; i know it aint gonna include any angle postulate

OpenStudy (anonymous):

Thanks